PLEASE USE RICCATI DE

T2-2 (20 Points): Find the G.S. of the DE: xy + y = 3x2 Prime denotes derivative WRT X. (Hint: guess a P.S. yı = Ax4)
Answer #1

Answer :

Consider the given differential equation

xy'+y=3x^2

Let y_{1} =Ax^{\alpha} be the particular solution

Then y_{1}^{'} =\alpha Ax^{\alpha-1}

By substituting y_{1} , \ y_{1}^{'} in the given differential equation

x\left ( {\alpha}Ax^{\alpha -1} \right )+Ax^{\alpha}=3x^2

\left ( {\alpha}Ax^{\alpha } \right )+Ax^{\alpha}=3x^2

\left ( {\alpha}+1\right )Ax^{\alpha}=3x^2

By equating , we get

{\alpha}=2 and A = 1

Thus the particular solution of the given differential equation is

y = x^{2}